### v1

# SORT3D: Spatial Object-centric Reasoning Toolbox for Zero-Shot 3D Grounding Using Large Language Models

Abstract
Interpreting object-referential language and grounding objects in 3D with spatial relations and attributes is essential for robots operating alongside humans. However, this task is often challenging due to the diversity of scenes, large number of fine-grained objects, and complex free-form nature of language references. Furthermore, in the 3D domain, obtaining large amounts of natural language training data is difficult. Thus, it is important for methods to learn from little data and zero-shot generalize to new environments. To address these challenges, we propose SORT3D, an approach that utilizes rich object attributes from 2D data and merges a heuristics-based spatial reasoning toolbox with the ability of large language models (LLMs) to perform sequential reasoning. Importantly, our method does not require text-to-3D data for training and can be applied zero-shot to unseen environments. We show that SORT3D achieves state-of-the-art performance on complex view-dependent grounding tasks on two benchmarks. We also implement the pipeline to run real-time on an autonomous vehicle and demonstrate that our approach can be used for object-goal navigation on previously unseen real-world environments. All source code for the system pipeline is publicly released.  [^1] https://github.com/nzantout/SORT3D
I INTRODUCTION
Figure 1: An example of our system’s workflow for using referential object grounding for downstream object-goal navigation. The agent uses the 2D image for fine-grained grounding in the presence of distractor objects.
As we progress toward generalizable robots operating in human-centered environments like homes and offices, it is crucial for these agents to interact through natural language and align visual observations with natural language references. This capability is essential for applications such as indoor navigation robots and robot caregiving assistants. Resolving natural language expressions referring to specific objects using spatial relations—the core challenge of 3D Visual Grounding (3DVG)—remains difficult despite being an easy task for humans. For example, understanding statements such as “the chair closest to the closet door”, is a task trivial for humans [1] but still challenging for robots. While humans are usually able to identify objects from referring expressions by filtering out irrelevant objects, reasoning about spatial relationships, and utilizing semantic object attributes, such tasks remain challenging for state-of-the-art (SOTA) methods due to several reasons. First, indoor environments often contain numerous objects belonging to fine-grained classes [2], and are often specific to different environments and homes. Second, grounding complex language references to a 3D scene is difficult even for SOTA large language models (LLMs), especially when the language contains view-dependent references (e.g. “facing the door”) or use unique ways of anchoring objects that require numerical reasoning (e.g. “the second chair from the window”). Third, to train learning-based methods on such grounding tasks, a large amount of annotated data aligning language references to a 3D scene is required, which the 3D domain lacks in comparison to the 2D vision-language domain [1].
While a number of existing works have developed end-to-end methods to tackle this task through training multi-modal alignment with large transformer models [3, 4, 5], these methods require large-scale annotated data, often overfitting to specific syntactic structure in training datasets while struggling to generalize to more complex utterances, resulting in only mediocre performance. More recently, numerous works have leveraged the reasoning abilities of LLMs for spatial grounding [6, 7, 8, 9, 10, 11, 12]. While some achieve strong results on structured, synthetic datasets [6], complex real-world natural language expressions such as “the tall recycling bin to the left if you are facing the door” remain challenging. These methods also often rely on careful fine-tuning or struggle with poor zero-shot performance. Additionally, such methods are not typically designed for system deployment, relying on heavily processed high fidelity reconstructions of 3D scenes [13, 14] and failing to account for constraints in model size, efficiency, and noise in the real-time semantic mapping process.
To this end, we propose SORT3D, a Spatial Object-centric Reasoning Toolbox for 3D Grounding Using LLMs, shown in Fig. 2, as a novel pipeline for 3D spatial reasoning tailored towards the downstream application of object-goal navigation [15]. SORT3D is modeled intuitively after human reasoning in object disambiguation, enabling data-efficient 3D grounding without requiring annotated 3D training data. Instead, we leverage SOTA 2D vision-language models (VLMs) to extract semantic object attributes important for distinguishing objects. To efficiently filter relevant objects in large scenes with hundreds of objects, we use an open-vocabulary object filtering module, similar to how humans focus their attention to relevant objects. For complex spatial references, we utilize large language models (LLMs) with chain-of-thought [16], prompting them to reason sequentially. However, recognizing LLMs’ limitations in mathematical reasoning and spatial understanding, we introduce a modular spatial reasoning toolbox with predefined elementary spatial relations. This delegates complex spatial reasoning to structured rule-based logic, ensuring greater accuracy and interpretability. As a result, our method only requires a single in-context example of the toolbox usage and no other training data.
We evaluate our method on standard 3D object referential grounding benchmarks, ReferIt3D [1] and VLA-3D [17], and show that it is competitive with SOTA performance on complex view-dependent statements while requiring less data to zero-shot generalize. We also deploy our full pipeline on a robot ground vehicle for real-time indoor navigation, demonstrating our methods ability to generalize to previously unseen, dynamic environments and run online alongside real-time perception and autonomy stacks.
II RELATED WORK
Object Referential Datasets
Using referring expressions to identify a target object, also defined as the 3DVG task, has been explored in a number of datasets such as ReferIt3D [1], ScanRefer [18], SceneVerse [19], and VLA-3D [17] in the 3D domain. Of these datasets, only the Nr3D subset of ReferIt3D contains human-generated natural language utterances, while the remaining datasets generate synthetic statements using template-based generation or LLMs to rephrase. While VLA-3D generates statements synthetically, the dataset also contains mentions of both static object attributes such as color, and relative size attributes. Compared to the template-generated datasets, the human-generated statements in Nr3D contain far more complex spatial grounding language, such as “Facing the three boxes, it is the box on the right that is on top of the larger blue box.”, which requires view-point grounding, and “The lamp to the left of the desk. NOT the lamp between the beds”, which contains negation, and “the chair closest to the metal appliance” which uses coarse object references. However this is the only dataset with such reasoning challenges, indicating a lack of data for complex 3DVG and motivating the need for 3DVG methods to be data efficient while learning complex spatial reasoning.
3D Referential Object Grounding
Grounding object references to 3D scenes has been explored more extensively by various methods since the introduction of benchmarks specific to the task, including BUTD-DETR [3], MVT [4], ViL3DRel [18], 3D-VisTA [5], NS3D [7], GPS trained on SceneVerse [19], and Transcrib3D [6]. While the majority of methods utilize fusion of multi-modal information in a large transformer model [3, 4, 18, 5, 19], more recent methods have leveraged neuro-symbolic frameworks [7] and LLMs [6] to effectively reason about spatial relations. Utilizing the reasoning capabilities of LLMs in the text domain, Transcrib3D achieves overall accuracies of 70.2% on Nr3D and 98.4% on Sr3D, outperforming all previous methods after skipping statements that do not contain the target object or were answered incorrectly by human annotators. To achieve this, Transcrib3D relies on iterative code generation and fine-tuning on incorrect answers with self-reasoned corrections. While significant progress in grounding accuracy has been made, the complex spatial reasoning in the Nr3D dataset, especially statements that involve egocentric view-point, continue to pose a challenge, especially for zero-shot methods [11, 20, 9]. We aim to address this gap with our method by combining structured heuristics with the sequential reasoning capabilities of LLMs.
Grounding Large Language Models in 3D
Recent efforts have also aimed to develop 3D foundation models capable of handling general 3D tasks. These models build upon pretrained LLMs, fine-tuning them on tokenized 3D data. By leveraging the strong reasoning capabilities of LLMs, these approaches bridge the gap between natural language understanding and 3D perception. One of the pioneering works in this space, 3D-LLM [21], adopts a 2D VLM as its backbone. To extract 3D features, it projects a given 3D input into multiple 2D images, obtains dense 2D features using ConceptFusion [22], and reconstructs 3D representations from these 2D embeddings. [8] improves upon the work to handle grounding to different viewpoints and generalizes to various 3D tasks with large-scale language-scene pretraining. Another significant contribution in this domain is LEO [23], which aims to create a unified 3D foundation model. LEO utilizes a pretrained LLM and applies LoRA [24] fine-tuning to adapt it for 3D-centric tasks. A key enabling factor for these methods is their reliance on pretrained LLMs and dense 2D data, demonstrating that grounding 3D understanding in well-established 2D and language-based representations is a powerful approach for advancing 3D reasoning. This factor subsequently inspires the design of our pipeline with the incorporation of 2D captions.
Vision-Language-Navigation with LLMs
A number of recent works have also leveraged the sequential reasoning capabilities of LLMs for navigation tasks. NavGPT [25] and NavCoT [26] use VLMs to generate text descriptions of viewpoints in the scene in 2D, then task the LLM with selecting the next action in an instruction-following task. Such methods demonstrate the effectiveness of leveraging 2D visual information to guide grounding and action selection in the 3D space. While our problem setting uses object-goal navigation as a downstream task, object-centric grounding is crucial for interpreting general navigation instructions as these instructions frequently use objects to anchor viewpoints or actions.
III METHODOLOGY
Figure 2: The full system diagram for the SORT3D framework
We formulate the 3DVG task as a precursor to downstream tasks such as object-goal navigation, instruction-following, or scene visual question answering (VQA). The input to the grounding pipeline consists of perception information from the scene and a free-form referring expression in natural language. The output is the ID of the target object referenced. Fig. 2 shows our proposed framework, which can be broken down into four components: perception of the 3D scene including a pipeline to incorporate 2D semantic information for each object, filtering for relevant objects based on the input utterance, LLM-based spatial reasoning to resolve the target object, and code generation for executing a downstream action. Each of these components is described in further detail below.
III-A Enhancing Object Perception with 2D Captions
Accurately understanding the attributes and affordances of 3D objects is an essential first step for referential grounding. Existing works [6], [4] use high-level information such as object bounding boxes and labels for this task, often acquired through 3D segmentation. While 3D segmentations provide useful object-centric information, they often fail to capture fine-grained attributes like color and shape, which is a key bottle-neck of 3D object grounding models[5]. However, large-scale training for accurate 3D perception is still a challenging problem due to the lack of data and task complexity. On the other hand, VLMs trained on vast amounts of 2D data have strong priors for object understanding in 2D. Notably, VQA models excel at captioning objects in a scene. Therefore, we leverage 2D VQA models to generate descriptions for 3D objects in the scene, providing richer visual details for grounding that are otherwise missed with pure 3D perception. This approach also mirrors how humans perceive and identify objects: narrowing down candidate objects through attributes and relations to resolve ambiguities. In our approach, this finer-grained inspection is achieved by captioning cropped object images, providing a more intuitive and precise form of object grounding.
A key decision to be made for captioning is what image to give the VQA model when multiple views of an object are present. We make this choice by using the viewpoint that has the highest CLIP similarity with the target label. We use the Qwen2-7B model as our VLM as we found it to work best to generate accurate and concise descriptions following our template. We query the VLM with the following prompt format: ”You are an AI model that describes the characteristics of a query object in an image. Describe the $<$object$>$ in this image, using properties like color, material, shape, affordances, and other meaningful attributes. Provide the response in this format: “The $<$object name$>$ is $<$color$>$, $<$material$>$, $<$shape$>$”. We release all object crops and captions as a supplement to the ScanNet [13] dataset along with our code. A sample of object crops and their corresponding captions are shown in Fig. 3.
Figure 3: Generated image crops and corresponding caption
III-B Filtering for Relevant Objects
Indoor environments such as homes can consist of hundreds of objects, few of which are relevant to a given language query or task. Thus, inspired by the ability for humans to filter out irrelevant objects and the success of past works [6, 7], we implement a filtering module that consists of: extracting referred objects along with their mentioned attributes from the input query, selecting a subset of relevant objects from the scene based on the mentioned properties, and only providing this subset of objects and properties to the downstream reasoning agent. We use the Mistral Large model [27] for these steps, filtering objects based on their text descriptions, to best leverage the ability of LLMs to process textual information.
III-C Spatial Reasoning Toolbox
With only the relevant objects extracted, the subset of objects and their captions are then fed to an LLM reasoning agent as a list of $n_{o}$ objects, and each object $o_{i}$ is represented by the list of attributes $\{id,name,caption,c_{x},c_{y},c_{z},size\}$, where $id$ is a unique integer identifier for the object, and $(c_{x},c_{y},c_{z})$ are discretized coordinates of the object center. The LLM is provided with a spatial reasoning toolbox and prompted to use chain-of-thought reasoning to pick out the referred object in the referential statement. We parse the selection as the navigation target for the real-world experiment (section III-D).
Directly prompting LLMs for object grounding often results in poor performance due to their limitations in spatial and mathematical reasoning: for example, when querying ”the nightstand to the left of the bed”, the LLM picks the one with the smallest $x$ value. As a result, we create a spatial reasoning toolbox consisting of heuristic search functions that find objects referred to by spatial relations and order them in likelihood of satisfying the relations. We then prompt the LLM to delegate object grounding to sequences of calls to the toolbox, which effectively maps operations that require complex spatial reasoning into sequential reasoning and set union/intersection problems. To specifically handle view-dependent queries, we leverage knowledge of free traversable space to anchor feasible viewpoints. During prompting of the LLM, only one in-context example of spatial reasoning with tool calls to demonstrate their usage is required. The modular architecture of our toolbox allows for easy expansion or modification of provided tools.
III-D Parsing for Action Execution
Finally, we tailor our pipeline for downstream navigation tasks by parsing LLM generated actions to either go near a given waypoint or go between two waypoints based on the target object selected. We first convert the target object location into a navigation waypoint. Given the center point of the selected target object, we calculate the closest point in free-space in the scene traversable by a robot agent and call on action functions with this waypoint.
IV EXPERIMENTAL SETUP
We evaluate our method on two 3D object-referential datasets, ReferIt3D [1] and VLA-3D [17]. Both datasets consist of utterances describing a target object in a Scannet [13] scene using spatial relations. In particular, ReferIt3D is split into Sr3D, which consists of synthetically generated utterances from five relation categories while Nr3D consists of natural language statements collected from humans, with unconstrained methods of describing target objects. Statements are categorized as “Easy”/“Hard” based on the number of “distractor” objects of the same class as the target object in the scene and also “View-Dependent”/“View-Independent”. VLA-3D consists entirely of template-generated statements that are view-independent but contains utterances for a diverse set of 3D scans and enforces every statement to contain a spatial relation. The set of spatial relations is expanded to include eight total relations, including ternary relations (e.g. “between”) and numerical relations (e.g. “second closest”). Utterances may also contain attributes such as color and size if needed to disambiguate the target object from distractors.
IV-A 3D Scene Grounding
We test our model on both ReferIt3D subsets and the subset of VLA-3D using Scannet scenes and compare to SOTA baselines. On each data subset, we evaluate our model on 200 sampled statements  [^2] We use a subset of the test set due to the cost of LLM evaluation on the full test set and the need for multiple runs to obtain variance statistics, sampled to match the distribution of Easy, Hard, View-Dependent, and View-Independent statements in the original ReferIt3D test dataset. We focus our comparison against Transcrib3d [6] which, to the best of our knowledge, is the best performing model on ReferIt3D to date. We note, however, that their evaluation skipped over statements in the test set that either do not explicitly mention the target object or human annotators were unable to guess correctly during the dataset generation. We provide results for our method on both cases, with these statements included and skipped, where “+skip” indicates the exclusion of such statements and subsequent simplification of the problem. For a fair comparison, we run Transcrib3d with the same two LLMs we use on the same test splits 333the GPT4 model used in their work is now a legacy model. We run their method with GPT4o instead.. For our methods, we conduct multiple trials on each data split to measure variance in LLMs, reported with standard deviation values on the grounding accuracy, which we note that other LLM-based methods do not report.
V RESULTS AND DISCUSSION
V-A 3D Scene Grounding
The grounding accuracy on ReferIt3D is shown in Tables I and II, and accuracy on VLA-3D is shown in Table III. We see that our method achieves higher accuracy with GPT-4o as the LLM backend and is on par with SOTA methods on View-Dependent statements in Nr3D and Hard statements in VLA-3D while requiring no data to train. We also note that the use of LLMs introduces variance between trials, affecting grounding accuracy up to 6%.
While supervised baselines such as ViL3DRel[28], 3D-VisTA [5], and SceneVerse [19] report slightly higher overall grounding accuracies, these methods are explicitly trained on ReferIt3D data. Similarly, while Transcrib3D [6] reports higher accuracies, it relies on a) removing statements that annotators incorrectly answered and b) iterative self-correction, where incorrect queries are fed back with oracle labels as few-shot examples for fine-tuning. While this approach improves accuracy on the benchmark, it arguably ”forces” the LLM to remember its prior mistakes within the dataset.
In contrast, our approach is purely zero-shot, requiring only one single example of how to use the spatial reasoning toolbox, which does not have to be from a particular dataset, and we employed no dataset-specific training or fine-tuning. Despite this, by mapping spatial reasoning into sequential reasoning, our spatial reasoning toolbox approach achieves performance comparable to SOTA supervised and fine-tuned methods on view-dependent statements in Nr3D, and surpasses other zero-shot methods on grounding accuracy, particularly on view-dependent statements. On Sr3D, SORT3D surpasses SOTA supervised training methods while achieving closer performance to fine-tuned methods. This demonstrates the effectiveness of our approach at handling spatial reasoning where viewpoint anchoring is required. We see that SORT3D is able to explainably resolve complex view-dependent relations with multiple anchors and complex semantic descriptions (Fig. 4-a), and are additionally able to explain model failure points by analyzing its chain of thought (Fig. 4-b).
On VLA-3D, our method surpasses other methods on Hard statements by a large margin. VLA-3D contains a large set of statements using size and color descriptions when referring to objects, which our method effectively utilizes through object filtering, 2D caption descriptions, and the spatial reasoning toolbox to identify target objects with multiple distractors in the scene. Transcrib3D [6], when not fine-tuned toward the dataset’s hard statements, failed to generalize well despite using the same LLM backbone. The results on VLA-3D strongly support the benefits of incorporating 2D captions, as they provide critical, fine-grained object-level attributes that refine the referential grounding.
Figure 4: Correct (a) and incorrect (b) grounding examples. Top left and bottom left respectively show correctly grounded view-independent and view-dependent statements. Top right and bottom right are two examples of model logic failing: in the top right image, the model picks out the desk closest to a window, but not near the whiteboard. In the bottom right, the model fails at pragmatics, picking out the rightmost pillow, instead of recognizing that the sentence implies choosing a pillow on the bed.
TABLE I: Performance on the Nr3D dataset from the ReferIt3D benchmark. Asterisks (*) indicate results reported from the paper directly. “+skip” indicates exclusion of statements answered incorrectly by human annotators. “View Dep.” and “View Ind.” stand for view-dependent and view-independent respectively.
Nr3D
Method Overall View Dep. View Ind.
Supervised Methods
NS3D* [7] 62.7 62.0 -
ViL3DRel* [28] 64.4 62.0 64.5
3D-VisTA* [5] 64.2 61.5 65.1
SceneVerse-GPS* [19] 64.9 56.9 67.9
Fine-Tuned Methods
Transcrib3D* [6] (GPT-4+skip) 70.2 60.1 75.4
Transcrib3D (GPT-4o) 58.5 54.0 60.6
Transcrib3D (Mistral) 58.5 52.4 61.3
Transcrib3D (GPT-4o+skip) 65.6 63.3 66.7
Transcrib3D (Mistral+skip) 63.8 57.1 66.7
Zero-Shot Methods
ZSVG3D [20] 39.0 36.8 40.0
VLM-Grounder [11] 48.0 45.8 49.4
CSVG [9] 59.2 53.0 62.5
Ours (Mistral) 56.5$\pm$0.5 57.6$\pm$3.6 56.2$\pm$2.3
Ours (GPT-4o) 60.5$\pm$1.0 56.6$\pm$3.3 62.3$\pm$2.4
Ours (GPT-4o+skip) 62.0$\pm$1.2 56.6$\pm$0.0 64.3$\pm$1.7
Ours (Mistral+skip) 61.6$\pm$0.3 59.4$\pm$0.9 62.6$\pm$0.9
TABLE II: Performance on the Sr3D dataset from the ReferIt3D benchmark. Asterisks (*) indicate results reported from the paper directly.
Sr3D
Method Overall View Dep. View Ind.
Supervised Methods
ViL3DRel* [28] 72.8 63.8 73.2
3D-VisTA* [5] 76.4 58.9 77.3
SceneVerse-GPS* [19] 77.5 62.8 78.2
Fine-Tuned Methods
Transcrib3D* [6] (GPT-4+skip) 98.4 98.2 98.4
Transcrib3D (GPT-4o) 96.5 88.9 96.9
Transcrib3D (Mistral) 96.0 77.8 96.9
Zero-Shot Methods
Ours (Mistral) 92.0$\pm$0.7 90.9$\pm$0.0 92.2$\pm$0.8
Ours (GPT-4o) 92.0$\pm$0.0 95.5$\pm$0.0 91.6$\pm$0.0
TABLE III: Grounding accuracy on VLA-3D test subset
VLA3D
Method Overall Easy Hard
Transcrib3D (Mistral) 70.5 76.25 47.5
Transcrib3D (GPT-4o) 77.5 82.5 57.5
Ours (Mistral) 69.0$\pm$0.7 70.0$\pm$0.8 65.0$\pm$0.0
Ours (GPT-4o) 71.8$\pm$1.8 71.0$\pm$1.3 75.0$\pm$3.5
V-B Ablation of Captioning Module
We evaluate the effect on grounding accuracy of adding open-vocabulary captions generated from 2D images of objects in the scene. We augment the Transcrib3D [6] baseline model with our captions for each object as additional information passed into the LLM reasoner. We hypothesize that these finer descriptions of object attributes will help the model in disambiguating objects when given free-form referential statements. We test this hypothesis through an ablation study using GPT-4o with both our method and Transcrib3D on Nr3D, shown in Table IV. In both Transcrib3D and our method, we observe consistent and significant improvements across all statements types. For our approach specifically, the addition of captions improves performance the most (11%) on view-dependent statements. These results demonstrate that understanding detailed object attributes is important for effective 3D grounding and leveraging 2D VLMs is an effective method for this. Many referential statements rely on subtle distinctions—such as color, shape, texture, or affordance—that traditional 3D models often miss.
V-C Real-World Validation
To validate SORT3D’s pipeline in the real-world, we implemented the full system on a differential-drive ground vehicle equipped with a 360-degree camera, a Velodyne LiDAR sensor, an onboard Intel NUC for low level autonomy, and a desktop with an RTX 4090. The system starts with an empty map and continuously integrates real-time RGB-D input to perform panoptic segmentation. This enables the robot to simultaneously build a map, detect objects, and assign open-vocabulary labels, allowing it to navigate and understand its environment dynamically. As the robot explores, it captures images of objects and our pipeline selects image crops and generates open-vocabulary captions, then responds to grounding language queries whenever they are typed in real-time, triggering the rest of the SORT3D pipeline. We validate our system in a previously unseen environment, demonstrating its ability to handle complex statements including compound statements and open-vocabulary object-centric navigation commands, as shown in Fig 6.
Figure 5: An overhead view of the real world kitchen environment and the differential-drive robot used in the experiment.
Figure 6: Real-world examples of SORT3D referential grounding and navigation on compound relational statements and multi-object navigation 3D. The top row shows the egocentric camera view of the robot and the bottom row shows the path taken in a map of the scene. The target objects are highlighted in green and the anchor objects blue.
VI LIMITATIONS AND FUTURE WORK
We acknowledge that limitations exist within in our approach. First, our system relies on internet access to call online APIs, which is a reasonable assumption for indoor environments like homes but may not hold for generalizable robot navigation systems. To address this limitation, the modular design of our pipeline allows for the LLMs to be easily replaced by smaller local models.
We also recognize that our evaluation set is limited. There are few existing benchmarks designed to rigorously test online 3D referential grounding with diverse, attribute-rich natural language. This constraint makes it challenging to fully assess the generalizability of our approach across different environments or while deployed on a system. Evaluating our method in a simulated environment with real-time interactions would provide deeper insights into its effectiveness and adaptability as a system.
TABLE IV: Grounding accuracy with and without captions on Nr3D
Nr3D
Method Overall Easy Hard View Dep. View Ind.
Transcrib3D (GPT-4o) 58.5 67.5 45.0 54.0 60.6
Transcrib3D (GPT-4o + captions) 61.0 ($\uparrow$ 4.3%) 70.8 ($\uparrow$ 4.9%) 46.3 ($\uparrow$ 2.9%) 55.4 ($\uparrow$ 2.6%) 63.5 ($\uparrow$ 4.8%)
Ours (GPT-4o) 54.5 59.2 47.5 50.7 56.2
Ours (GPT-4o + captions) 60.5 ($\uparrow$ 11.0%) 64.2 ($\uparrow$ 8.4%) 55.0 ($\uparrow$ 5.3%) 56.6 ($\uparrow$ 11.6%) 62.3 ($\uparrow$ 10.9%)
VII CONCLUSION
We introduce SORT3D, a robust, data-efficient, and deployable method for 3D referential grounding with complex view-dependent reasoning. Our framework combines 2D visual features, LLM-based filtering, and a heuristics-driven spatial reasoning toolbox to enhance LLMs’ ability to handle spatial and mathematical reasoning in 3D scenes. SORT3D achieves results that are competitive to SOTA methods on view-dependent and attribute-based statements while outperforming other fully zero-shot methods. We further demonstrate its performance and generalization by deploying the pipeline on a robotic system for real-time indoor navigation.

### v2

@@
 Importantly, our method does not require text-to-3D data for training and can be applied zero-shot to unseen environments.
-We show that SORT3D achieves state-of-the-art performance on complex view-dependent grounding tasks on two benchmarks.
-We also implement the pipeline to run real-time on an autonomous vehicle and demonstrate that our approach can be used for object-goal navigation on previously unseen real-world environments.
+We show that SORT3D achieves state-of-the-art zero-shot performance on complex view-dependent grounding tasks on two benchmarks.
+We also implement the pipeline to run real-time on two autonomous vehicles and demonstrate that our approach can be used for object-goal navigation on previously unseen real-world environments.
 All source code for the system pipeline is publicly released.
@@
 As we progress toward generalizable robots operating in human-centered environments like homes and offices, it is crucial for these agents to interact through natural language and align visual observations with natural language references.
-This capability is essential for applications such as indoor navigation robots and robot caregiving assistants.
-Resolving natural language expressions referring to specific objects using spatial relations—the core challenge of 3D Visual Grounding (3DVG)—remains difficult despite being an easy task for humans.
+This capability is essential for applications such as robot caregivers and indoor assistants.
+Resolving natural language expressions referring to specific objects using semantic object attributes and inter-object spatial relations—the core challenge of 3D referential grounding—remains difficult despite being an intuitive task for humans.
 For example, understanding statements such as “the chair closest to the closet door”, is a task trivial for humans [1] but still challenging for robots.
 While humans are usually able to identify objects from referring expressions by filtering out irrelevant objects, reasoning about spatial relationships, and utilizing semantic object attributes, such tasks remain challenging for state-of-the-art (SOTA) methods due to several reasons.
-First, indoor environments often contain numerous objects belonging to fine-grained classes [2], and are often specific to different environments and homes.
-Second, grounding complex language references to a 3D scene is difficult even for SOTA large language models (LLMs), especially when the language contains view-dependent references (e.g.
-“facing the door”) or use unique ways of anchoring objects that require numerical reasoning (e.g.
-“the second chair from the window”).
-Third, to train learning-based methods on such grounding tasks, a large amount of annotated data aligning language references to a 3D scene is required, which the 3D domain lacks in comparison to the 2D vision-language domain [1].
-While a number of existing works have developed end-to-end methods to tackle this task through training multi-modal alignment with large transformer models [3, 4, 5], these methods require large-scale annotated data, often overfitting to specific syntactic structure in training datasets while struggling to generalize to more complex utterances, resulting in only mediocre performance.
-More recently, numerous works have leveraged the reasoning abilities of LLMs for spatial grounding [6, 7, 8, 9, 10, 11, 12].
-While some achieve strong results on structured, synthetic datasets [6], complex real-world natural language expressions such as “the tall recycling bin to the left if you are facing the door” remain challenging.
-These methods also often rely on careful fine-tuning or struggle with poor zero-shot performance.
-Additionally, such methods are not typically designed for system deployment, relying on heavily processed high fidelity reconstructions of 3D scenes [13, 14] and failing to account for constraints in model size, efficiency, and noise in the real-time semantic mapping process.
-To this end, we propose SORT3D, a Spatial Object-centric Reasoning Toolbox for 3D Grounding Using LLMs, shown in Fig.
-2, as a novel pipeline for 3D spatial reasoning tailored towards the downstream application of object-goal navigation [15].
-SORT3D is modeled intuitively after human reasoning in object disambiguation, enabling data-efficient 3D grounding without requiring annotated 3D training data.
-Instead, we leverage SOTA 2D vision-language models (VLMs) to extract semantic object attributes important for distinguishing objects.
-To efficiently filter relevant objects in large scenes with hundreds of objects, we use an open-vocabulary object filtering module, similar to how humans focus their attention to relevant objects.
-For complex spatial references, we utilize large language models (LLMs) with chain-of-thought [16], prompting them to reason sequentially.
-However, recognizing LLMs’ limitations in mathematical reasoning and spatial understanding, we introduce a modular spatial reasoning toolbox with predefined elementary spatial relations.
-This delegates complex spatial reasoning to structured rule-based logic, ensuring greater accuracy and interpretability.
+First, indoor environments often contain numerous objects belonging to fine-grained classes [2], with distributions that vary widely across different homes and environments.
+Second, training end-to-end learning-based methods on 3D referential grounding requires a large amount of annotated data aligning language references to a 3D scene, which the 3D domain lacks in comparison to the 2D vision-language domain [1].
+While a number of existing works have developed end-to-end methods to tackle this task through training multi-modal alignment with large transformer models [3, 4, 5], these methods require large-scale annotated data, often overfitting to specific syntactic structure in training datasets while struggling to generalize to more complex utterances, resulting in mediocre performance.
+More recently, numerous works have leveraged the reasoning capabilities and rich language semantics of large language models (LLMs) for 3D referential grounding [6, 7, 8, 9, 10, 11, 12].
+While some achieve strong results on benchmark datasets [6], complex real-world natural language expressions that employ both semantic attributes and spatial relations such as “the tall recycling bin to the left if you are facing the door” remain challenging.
+Many LLM-based approaches either struggle with poor zero-shot performance or rely on careful fine-tuning or heavy prompt engineering tailored towards benchmark datasets.
+Consequently, such methods are not typically designed for system deployment, relying on high fidelity reconstructed meshes of 3D scenes [13, 14] and failing to account for constraints in model size, efficiency, and noise in the real-time semantic mapping process.
+To this end, we propose SORT3D, a Spatial Object-centric Reasoning Toolbox for 3D Grounding Using LLMs, shown in Figure 2, as a novel pipeline for 3D spatial reasoning tailored towards the downstream application of object-goal navigation [15].
+SORT3D is modeled intuitively after human reasoning in object disambiguation, enabling data-efficient 3D referential grounding without requiring annotated 3D training data.
+To achieve this, we decompose the task into a three-stage approach.
+After obtaining object names and bounding boxes using an instance-level semantic mapping module, we leverage SOTA 2D vision-language models (VLMs) to extract semantic object attributes important for distinguishing objects.
+We then use an open-vocabulary object filtering module to efficiently filter relevant objects in large scenes with hundreds of objects, similar to how humans focus their attention to relevant objects.
+We finally leverage the strong semantic language and sequential reasoning priors of LLMs using chain-of-thought prompting [16] to parse a complex referential statement into a series of function calls to our spatial reasoning toolbox, containing search functions that return the IDs of objects that satisfy heuristically defined elementary spatial relations.
+This delegates complex relative spatial reasoning to structured rule-based logic, ensuring greater accuracy and interpretability.
 As a result, our method only requires a single in-context example of the toolbox usage and no other training data.
-We evaluate our method on standard 3D object referential grounding benchmarks, ReferIt3D [1] and VLA-3D [17], and show that it is competitive with SOTA performance on complex view-dependent statements while requiring less data to zero-shot generalize.
-We also deploy our full pipeline on a robot ground vehicle for real-time indoor navigation, demonstrating our methods ability to generalize to previously unseen, dynamic environments and run online alongside real-time perception and autonomy stacks.
+We evaluate our method on standard 3D object referential grounding benchmarks, ReferIt3D [1] and IRef-VLA [17], and demonstrate performance competitive with SOTA on complex view-dependent statements while requiring less data to zero-shot generalize.
+We also deploy our full pipeline on two robotic ground vehicles for real-time indoor navigation, demonstrating our method’s ability to further generalize to previously unseen environments and run online alongside real-time perception and autonomy stacks.
 II Related Work
-Object Referential Datasets
-Using referring expressions to identify a target object, also defined as the 3DVG task, has been explored in a number of datasets such as ReferIt3D [1], ScanRefer [18], SceneVerse [19], and VLA-3D [17] in the 3D domain.
-Of these datasets, only the Nr3D subset of ReferIt3D contains human-generated natural language utterances, while the remaining datasets generate synthetic statements using template-based generation or LLMs to rephrase.
-While VLA-3D generates statements synthetically, the dataset also contains mentions of both static object attributes such as color, and relative size attributes.
-Compared to the template-generated datasets, the human-generated statements in Nr3D contain far more complex spatial grounding language, such as “Facing the three boxes, it is the box on the right that is on top of the larger blue box.”, which requires view-point grounding, and “The lamp to the left of the desk.
-NOT the lamp between the beds”, which contains negation, and “the chair closest to the metal appliance” which uses coarse object references.
-However this is the only dataset with such reasoning challenges, indicating a lack of data for complex 3DVG and motivating the need for 3DVG methods to be data efficient while learning complex spatial reasoning.
-3D Referential Object Grounding
-Grounding object references to 3D scenes has been explored more extensively by various methods since the introduction of benchmarks specific to the task, including BUTD-DETR [3], MVT [4], ViL3DRel [18], 3D-VisTA [5], NS3D [7], GPS trained on SceneVerse [19], and Transcrib3D [6].
-While the majority of methods utilize fusion of multi-modal information in a large transformer model [3, 4, 18, 5, 19], more recent methods have leveraged neuro-symbolic frameworks [7] and LLMs [6] to effectively reason about spatial relations.
-Utilizing the reasoning capabilities of LLMs in the text domain, Transcrib3D achieves overall accuracies of 70.2% on Nr3D and 98.4% on Sr3D, outperforming all previous methods after skipping statements that do not contain the target object or were answered incorrectly by human annotators.
-To achieve this, Transcrib3D relies on iterative code generation and fine-tuning on incorrect answers with self-reasoned corrections.
-While significant progress in grounding accuracy has been made, the complex spatial reasoning in the Nr3D dataset, especially statements that involve egocentric view-point, continue to pose a challenge, especially for zero-shot methods [11, 20, 9].
-We aim to address this gap with our method by combining structured heuristics with the sequential reasoning capabilities of LLMs.
+Referential Object Grounding Datasets
+Using referring expressions to identify a target object in a 3D scene, defined as the 3D referential grounding task, has been explored in a number of datasets such as ReferIt3D [1], ScanRefer [18], SceneVerse [19], and IRef-VLA [17].
+Of these datasets, only the Nr3D subset of ReferIt3D and ScanRefer contain human-generated natural language utterances, while the remaining datasets generate synthetic statements using template-based generation and/or LLMs to rephrase.
+Synthetic datasets only employ basic spatial relations like ”lamp by the desk” in referential statements, although IRef-VLA adds the usage basic semantic object attributes—color and size—to create unambiguous references.
+Compared to the template-generated datasets, the human-generated statements in Nr3D contain far more complex spatial grounding language, such as “Facing the three boxes, it is the box on the right that is on top of the larger blue box.”, which requires view-point grounding, “The lamp to the left of the desk.
+NOT the lamp between the beds”, which contains negation, and “the chair closest to the metal appliance” which uses coarse object references and semantic attribute.
+This complexity, coupled with the small scale of human-generated 3D referential language data, motivates the need for 3D referential grounding methods to leverage strong semantic language priors from other sources (like LLMs), and use a more structured approach for grounding.
+3D Referential Object Grounding Baselines
+Grounding object references to 3D scenes has been explored more extensively by various methods since the introduction of benchmarks specific to the task.
+Approaches to this task include end-to-end models like BUTD-DETR [3], MVT [4], ViL3DRel [20], 3D-VisTA [5], and GPS trained on SceneVerse [19], which fuse multi-modal information in large transformer models and are trained and fine-tuned directly on referential grounding benchmarks.
+More recent methods decompose the task, leveraging neuro-symbolic frameworks like NS3D [7], and LLMs and VLMs in a zero-shot manner to effectively reason about spatial relations like ZSVG3D [21], VLM-Grounder [11], CSVG [9], and Transcrib3D [6].
+Utilizing the reasoning capabilities of LLMs in the text domain, Transcrib3D achieves overall accuracies of 70.2% and 98.4% respectively on subsets of Nr3D and Sr3D, outperforming all previous methods.
+To achieve this, Transcrib3D represents a scene as a list of object names, colors, and positions, and relies on iterative code generation and benchmark-specific prompt engineering giving the LLM guiding principles for grounding.
+Transcrib3D additionally fine-tunes smaller LLMs on incorrect answers with corrections self-reasoned by larger LLMs.
+While significant progress in grounding accuracy has been made, the complex spatial reasoning required for the Nr3D dataset, especially with statements that involve egocentric viewpoints and utilize object semantics, continues to pose a challenge, especially for zero-shot methods [11, 21, 9, 6].
+We aim to address this gap with our method by combining structured heuristics and vision foundation models with the sequential reasoning capabilities of LLMs.
 Grounding Large Language Models in 3D
 Recent efforts have also aimed to develop 3D foundation models capable of handling general 3D tasks.
-These models build upon pretrained LLMs, fine-tuning them on tokenized 3D data.
-By leveraging the strong reasoning capabilities of LLMs, these approaches bridge the gap between natural language understanding and 3D perception.
-One of the pioneering works in this space, 3D-LLM [21], adopts a 2D VLM as its backbone.
-To extract 3D features, it projects a given 3D input into multiple 2D images, obtains dense 2D features using ConceptFusion [22], and reconstructs 3D representations from these 2D embeddings.
+These models leverage the strong reasoning capabilities of pretrained LLMs, achieving strong performance on 3D scene understanding tasks by fine-tuning the LLMs on tokenized 3D data.
+One of the pioneering works in this space, 3D-LLM [22], distills rich 2D foundation model features and a 2D VLM backbone to improve performance, mitigating the lack of 3D data.
 [8] improves upon the work to handle grounding to different viewpoints and generalizes to various 3D tasks with large-scale language-scene pretraining.
-Another significant contribution in this domain is LEO [23], which aims to create a unified 3D foundation model.
-LEO utilizes a pretrained LLM and applies LoRA [24] fine-tuning to adapt it for 3D-centric tasks.
-A key enabling factor for these methods is their reliance on pretrained LLMs and dense 2D data, demonstrating that grounding 3D understanding in well-established 2D and language-based representations is a powerful approach for advancing 3D reasoning.
+The efficacy of 2D foundation model features in these approaches demonstrates that grounding 3D understanding in well-established 2D and language-based representations is a powerful approach for advancing 3D reasoning.
 This factor subsequently inspires the design of our pipeline with the incorporation of 2D captions.
-Vision-Language-Navigation with LLMs
+Vision-and-Language Navigation with LLMs
 A number of recent works have also leveraged the sequential reasoning capabilities of LLMs for navigation tasks.
@@
 Such methods demonstrate the effectiveness of leveraging 2D visual information to guide grounding and action selection in the 3D space.
-While our problem setting uses object-goal navigation as a downstream task, object-centric grounding is crucial for interpreting general navigation instructions as these instructions frequently use objects to anchor viewpoints or actions.
+However, within a collaborative setting between a human and a robotic agent, the human would more commonly use single referential statements like ”fetch the tv remote on the cabinet” and assume the robot has full knowledge the scene rather than describing a full trajectory towards an object.
+For building a 3D referential grounding-based VLN system with object-goal navigation as the basic downstream task, text descriptions of landmarks must also be combined with a structured map representing the scene.
 III Methodology
 Figure 2: The full system diagram for the SORT3D framework
-We formulate the 3DVG task as a precursor to downstream tasks such as object-goal navigation, instruction-following, or scene visual question answering (VQA).
+In a human-agent collaborative setting, humans commonly refer to objects in the scene by relating them to other objects using commands like ”grab the red mug in the top left cupboard over the sink”.
+Finding the correct object being referenced from the utterance is the task of 3D referential grounding, which is fundamental for the deployment of a practical VLN system.
+3D referential grounding additionally acts as a precursor to downstream tasks such as object-goal navigation, multi-action instruction-following, and scene visual question answering (VQA).
+We therefore present SORT3D, a zero-shot pipeline for 3D referential grounding, which decomposes the task into multiple subtasks, leverages foundation models to obtain robust zero-shot performance, and targets downstream mobile robot navigation in a real-world environment.
 The input to the grounding pipeline consists of perception information from the scene and a free-form referring expression in natural language.
 The output is the ID of the target object referenced.
-Fig.
-2 shows our proposed framework, which can be broken down into four components: perception of the 3D scene including a pipeline to incorporate 2D semantic information for each object, filtering for relevant objects based on the input utterance, LLM-based spatial reasoning to resolve the target object, and code generation for executing a downstream action.
+Figure 2 shows our proposed framework, which can be broken down into four components: (i) an instance-level semantic mapping system to obtain 3D bounding boxes for real-world deployment, (ii) a captioning pipeline to incorporate rich 2D semantic information for each object, (iii) filtering for relevant objects based on the input utterance, and (iv) LLM-based reasoning augmented with a spatial reasoning toolbox to resolve the target object, followed by code generation for executing a downstream action.
 Each of these components is described in further detail below.
-III-A Enhancing Object Perception with 2D Captions
+III-A Instance-level Semantic Mapping
+For our real-world experiments, we use an object instance-level semantic mapping module running in real-time to obtain the 3D bounding boxes to be input into the LLM and the spatial reasoning toolbox.
+This component is the only pipeine change required for our method to be deployed in the real-world.
+Our mobile robot perception setup for real-world experiments consists of a 360 camera and a 3D LiDAR (section V-C contains further hardware details).
+We initially perform object detection in 2D using open-vocabulary object detection [25] and instance segmentation [26] models.
+We then project the registered LiDAR point clouds onto the semantic images and associate each point with its corresponding pixel semantic ID.
+As the robot moves and produces new observations, we associate per-frame object instance pointclouds using a 2D tracking module and 3D proximity priors, followed by filtering steps to obtain 3D instance pointclouds [^2] For further details on the semantic mapping module, see https://github.com/gfchen01/semantic_mapping_with_360_camera_and_3d_lidar.
+The usage of open-vocabulary 2D foundation models allows our semantic mapping module to generalize to new environments as we show in our real-world experiments (section V-C).
+For our results on the ReferIt3D benchmarks, we simply use ground truth bounding boxes and instance segmentations.
+III-B Enhancing Object Perception with 2D Captions
 Accurately understanding the attributes and affordances of 3D objects is an essential first step for referential grounding.
 Existing works [6], [4] use high-level information such as object bounding boxes and labels for this task, often acquired through 3D segmentation.
-While 3D segmentations provide useful object-centric information, they often fail to capture fine-grained attributes like color and shape, which is a key bottle-neck of 3D object grounding models[5].
-However, large-scale training for accurate 3D perception is still a challenging problem due to the lack of data and task complexity.
+While 3D segmentation provides useful object-centric information, it often fails to capture fine-grained attributes like color and shape, which is a key bottleneck of 3D object grounding models [5].
+Nonetheless, large-scale training for accurate 3D perception is still a challenging problem due to the lack of data and task complexity.
 On the other hand, VLMs trained on vast amounts of 2D data have strong priors for object understanding in 2D.
@@
 This approach also mirrors how humans perceive and identify objects: narrowing down candidate objects through attributes and relations to resolve ambiguities.
-In our approach, this finer-grained inspection is achieved by captioning cropped object images, providing a more intuitive and precise form of object grounding.
+In our approach, this finer-grained inspection is achieved by generating captions for each object from cropped object images, providing a more intuitive and precise form of object grounding.
 A key decision to be made for captioning is what image to give the VQA model when multiple views of an object are present.
-We make this choice by using the viewpoint that has the highest CLIP similarity with the target label.
-We use the Qwen2-7B model as our VLM as we found it to work best to generate accurate and concise descriptions following our template.
-We query the VLM with the following prompt format: ”You are an AI model that describes the characteristics of a query object in an image.
-Describe the $<$object$>$ in this image, using properties like color, material, shape, affordances, and other meaningful attributes.
-Provide the response in this format: “The $<$object name$>$ is $<$color$>$, $<$material$>$, $<$shape$>$”.
+We make this choice by using the viewpoint that has the highest CLIP similarity with the target label, which is obtained from ground truth for the ReferIt3D benchmarks and from the semantic mapping module for the real-world experiments.
+We use Qwen2-VL-7B [27] as our VLM as we found it to perform best in generating accurate and concise descriptions following our template, and the quantized version of Qwen2.5-VL-Instruct-3B for system deployment due to memory constraints.
+We query the VLM with the following prompt format: ‘‘You are an AI model that describes the characteristics of a query object in an image.
+Describe the <<object>> in this image, using properties like color, material, shape, affordances, and other meaningful attributes.
+Provide the response in this format: ‘The <<object name>> is <<color>>, <<material>>, <<shape>>’".
 We release all object crops and captions as a supplement to the ScanNet [13] dataset along with our code.
-A sample of object crops and their corresponding captions are shown in Fig.
-3.
+A sample of object crops and their corresponding captions are shown in Figure 3.
 Figure 3: Generated image crops and corresponding caption
-III-B Filtering for Relevant Objects
-Indoor environments such as homes can consist of hundreds of objects, few of which are relevant to a given language query or task.
-Thus, inspired by the ability for humans to filter out irrelevant objects and the success of past works [6, 7], we implement a filtering module that consists of: extracting referred objects along with their mentioned attributes from the input query, selecting a subset of relevant objects from the scene based on the mentioned properties, and only providing this subset of objects and properties to the downstream reasoning agent.
-We use the Mistral Large model [27] for these steps, filtering objects based on their text descriptions, to best leverage the ability of LLMs to process textual information.
-III-C Spatial Reasoning Toolbox
-With only the relevant objects extracted, the subset of objects and their captions are then fed to an LLM reasoning agent as a list of $n_{o}$ objects, and each object $o_{i}$ is represented by the list of attributes $\{id,name,caption,c_{x},c_{y},c_{z},size\}$, where $id$ is a unique integer identifier for the object, and $(c_{x},c_{y},c_{z})$ are discretized coordinates of the object center.
-The LLM is provided with a spatial reasoning toolbox and prompted to use chain-of-thought reasoning to pick out the referred object in the referential statement.
-We parse the selection as the navigation target for the real-world experiment (section III-D).
-Directly prompting LLMs for object grounding often results in poor performance due to their limitations in spatial and mathematical reasoning: for example, when querying ”the nightstand to the left of the bed”, the LLM picks the one with the smallest $x$ value.
-As a result, we create a spatial reasoning toolbox consisting of heuristic search functions that find objects referred to by spatial relations and order them in likelihood of satisfying the relations.
-We then prompt the LLM to delegate object grounding to sequences of calls to the toolbox, which effectively maps operations that require complex spatial reasoning into sequential reasoning and set union/intersection problems.
-To specifically handle view-dependent queries, we leverage knowledge of free traversable space to anchor feasible viewpoints.
-During prompting of the LLM, only one in-context example of spatial reasoning with tool calls to demonstrate their usage is required.
-The modular architecture of our toolbox allows for easy expansion or modification of provided tools.
-III-D Parsing for Action Execution
-Finally, we tailor our pipeline for downstream navigation tasks by parsing LLM generated actions to either go near a given waypoint or go between two waypoints based on the target object selected.
-We first convert the target object location into a navigation waypoint.
-Given the center point of the selected target object, we calculate the closest point in free-space in the scene traversable by a robot agent and call on action functions with this waypoint.
+III-C Filtering for Relevant Objects
+Indoor environments such as homes can consist of hundreds of objects, of which only a few are relevant to a given language query or task.
+Thus, inspired by the ability for humans to filter out irrelevant objects and the success of past works [6, 7], we implement an LLM-based filtering module consisting of two queries.
+Given an input command like “The nightstand to the right of the bed”, the first query extracts object nouns and modifiers (e.g.
+nightstand and bed), and the second query returns the relevant object instances from the list of IDs and names provided by the perception module (in this case returning all nightstands and beds in the scene).
+We use Mistral Large 2 [28] for these steps, filtering objects based on their text descriptions, to best leverage the ability of LLMs to process textual information.
+III-D Spatial Reasoning Toolbox
+With only the relevant objects extracted, the subset of objects and their captions are then fed to an LLM reasoning agent as a list of $n_{o}$ objects, where each object $o_{i}$ is represented by the list of attributes: $\{\texttt{id},\texttt{name},\texttt{caption},c_{x},c_{y},c_{z},\texttt{size}\}$, where id is a unique integer identifier for the object, $(c_{x},c_{y},c_{z})$ are discretized coordinates of the object center, and size is the area of the largest face.
+However, directly prompting off-the-shelf LLMs for 3D referential grounding with this scene representation results in poor performance due to their limitations in spatial and mathematical reasoning: for example, when querying ”the nightstand to the left of the bed”, the LLM picks the one with the smallest $x$ value.
+As a result, we abstract spatial reasoning away from the LLM by creating a spatial reasoning toolbox consisting of heuristic search functions that find objects referred to by a set of elementary spatial relations.
+The key idea is that when referring to an object using relative spatial relations, only a fixed set of relations are needed.
+Thus, defining this set of heuristic functions is sufficient for handling diverse inter-object referential scenarios.
+The search functions and their arguments are listed in Table I.
+Spatial Search Functions
+find_near(target, anchor)*
+find_between(target, anchor1, anchor2)*
+find_above(target, anchor)*
+find_below(target, anchor)*
+order_bottom_to_top(targets)*
+order_smallest_to_largest(targets)*
+find_objects_near_room_corner(targets)*
+find_left(target, anchor)
+find_right(target, anchor)
+order_left_to_right(target, anchor)
+TABLE I: Heuristic search functions in the spatial toolbox. View-independent functions are marked with an asterisk (*), and view-dependent functions are marked with a dagger (†).
+The LLM is prompted with an in-context example to decompose a referential statement into a series of search calls and choose a single object ID from the returned lists of IDs.
+For example, given the query “Find the computer near the desk with a printer on it”, the LLM first calls find_below(desk, printer) returning [2], then find_near(computer, 2) returning [3, 4], finally picking the computer with ID 3.
+We note a few implementation details regarding the search functions:
+  1. 1.
+    For statements employing view-dependent relationships, like left and right, we tackle the case where no specific observer viewpoint is given (common in Nr3D), and assume the relationship is unambiguous from any feasible viewing direction.
+We therefore determinine left/right direction from the signed angle formed by the nearest point in free space and the target and anchor’s centroids.
+  2. 2.
+    We use the area of the largest face in order_smallest_to_largest, which is more intuitive than volume for flat objects.
+  3. 3.
+    All search functions have access to the full output of the perception module including object names and 3D bounding boxes, which is partially hidden from the LLM.
+This allows us to develop more complex logic for search functions without affecting LLM reasoning, which is an advantage of our approach.
+Our accompanying repository contains further implementation details on the spatial search functions.
+III-E Parsing for Action Execution
+For our benchmark results, the LLM is prompted to only output a single object ID denoting the chosen object.
+For our real-world deployed pipeline, the LLM may output a series of either go_near or go_between function calls which take object IDs as arguments and generate waypoints in the scene which are sent to an obstacle-avoidant downstream planner for sequential navigation.
+go_near(target) places a waypoint at the closest point in traversable space to the target’s center, while go_between(target1, target2) places one near the midpoint between the two targets.
 IV Experimental Setup
-We evaluate our method on two 3D object-referential datasets, ReferIt3D [1] and VLA-3D [17].
+We quantitatively evaluate our method on two 3D object-referential datasets, ReferIt3D [1] and IRef-VLA [17].
 Both datasets consist of utterances describing a target object in a ScanNet [13] scene using spatial relations.
@@
 Statements are categorized as “Easy”/“Hard” based on the number of “distractor” objects of the same class as the target object in the scene and also “View-Dependent”/“View-Independent”.
-VLA-3D consists entirely of template-generated statements that are view-independent but contains utterances for a diverse set of 3D scans and enforces every statement to contain a spatial relation.
+IRef-VLA consists entirely of template-generated statements that are view-independent but contains utterances for a diverse set of 3D scans and enforces every statement to contain a spatial relation.
 The set of spatial relations is expanded to include eight total relations, including ternary relations (e.g.
@@
 Utterances may also contain attributes such as color and size if needed to disambiguate the target object from distractors.
-IV-A 3D Scene Grounding
-We test our model on both ReferIt3D subsets and the subset of VLA-3D using Scannet scenes and compare to SOTA baselines.
+Additionally, we deploy SORT3D on two ground vehicles, and validate the system’s generalizability by testing navigation commands containing references to spatial relations, references to object attributes, and implicit or indirect requests in three previously unseen indoor environments.
+IV-A Referential Grounding on Benchmark Datasets
+We test our model on both ReferIt3D subsets and the subset of IRef-VLA using ScanNet scenes and compare to SOTA baselines.
 On each data subset, we evaluate our model on 200 sampled statements [^3] We use a subset of the test set due to the cost of LLM evaluation on the full test set and the need for multiple runs to obtain variance statistics, sampled to match the distribution of Easy, Hard, View-Dependent, and View-Independent statements in the original ReferIt3D test dataset.
 We focus our comparison against Transcrib3D [6] which, to the best of our knowledge, is the best performing model on ReferIt3D to date.
-We note, however, that their evaluation skipped over statements in the test set that either do not explicitly mention the target object or human annotators were unable to guess correctly during the dataset generation.
-We provide results for our method on both cases, with these statements included and skipped, where “+skip” indicates the exclusion of such statements and subsequent simplification of the problem.
-For a fair comparison, we run Transcrib3d with the same two LLMs we use on the same test splits 333the GPT4 model used in their work is now a legacy model.
+For a fair comparison, we run Transcrib3D with the same two LLMs we use on the same test splits444the GPT4 model used in their work is now a legacy model.
 We run their method with GPT4o instead..
@@
 V Results and Discussion
-V-A 3D Scene Grounding
-The grounding accuracy on ReferIt3D is shown in Tables I and II, and accuracy on VLA-3D is shown in Table III.
-We see that our method achieves higher accuracy with GPT-4o as the LLM backend and is on par with SOTA methods on View-Dependent statements in Nr3D and Hard statements in VLA-3D while requiring no data to train.
+V-A Referential Grounding on Benchmark Datasets
+The grounding accuracy on ReferIt3D is shown in Tables II and III, and accuracy on IRef-VLA is shown in Table IV.
+We see that our method achieves higher accuracy with GPT-4o as the LLM backend and is on par with SOTA methods on View-Dependent statements in Nr3D and Hard statements in IRef-VLA while requiring no data to train.
 We also note that the use of LLMs introduces variance between trials, affecting grounding accuracy up to 6%.
 While supervised baselines such as ViL3DRel [20], 3D-VisTA [5], and SceneVerse [19] report slightly higher overall grounding accuracies, these methods are explicitly trained on ReferIt3D data.
-Similarly, while Transcrib3D [6] reports higher accuracies, it relies on a) removing statements that annotators incorrectly answered and b) iterative self-correction, where incorrect queries are fed back with oracle labels as few-shot examples for fine-tuning.
-While this approach improves accuracy on the benchmark, it arguably ”forces” the LLM to remember its prior mistakes within the dataset.
+Similarly, while Transcrib3D reports higher accuracies, it relies on guiding principles [6] that are tailored to the language used in Nr3D and Sr3D, which improve performance on those benchmarks.
 In contrast, our approach is purely zero-shot, requiring only one single example of how to use the spatial reasoning toolbox, which does not have to be from a particular dataset, and we employed no dataset-specific training or fine-tuning.
-Despite this, by mapping spatial reasoning into sequential reasoning, our spatial reasoning toolbox approach achieves performance comparable to SOTA supervised and fine-tuned methods on view-dependent statements in Nr3D, and surpasses other zero-shot methods on grounding accuracy, particularly on view-dependent statements.
-On Sr3D, SORT3D surpasses SOTA supervised training methods while achieving closer performance to fine-tuned methods.
+Despite this, by leveraging foundation models to obtain object semantic attributes and mapping spatial reasoning into sequential reasoning, our spatial reasoning toolbox approach achieves overall performance comparable to SOTA supervised and fine-tuned methods on Nr3D, and performance on view-dependent statements on par with Transcrib3D.
+On Sr3D, SORT3D surpasses SOTA supervised training methods and achieves close overall performance to Transcrib3D while surpassing it in view-dependent accuracy.
 This demonstrates the effectiveness of our approach at handling spatial reasoning where viewpoint anchoring is required.
-We see that SORT3D is able to explainably resolve complex view-dependent relations with multiple anchors and complex semantic descriptions (Fig.
-4-a), and are additionally able to explain model failure points by analyzing its chain of thought (Fig.
-4-b).
-On VLA-3D, our method surpasses other methods on Hard statements by a large margin.
-VLA-3D contains a large set of statements using size and color descriptions when referring to objects, which our method effectively utilizes through object filtering, 2D caption descriptions, and the spatial reasoning toolbox to identify target objects with multiple distractors in the scene.
-Transcrib3D [6], when not fine-tuned toward the dataset’s hard statements, failed to generalize well despite using the same LLM backbone.
-The results on VLA-3D strongly support the benefits of incorporating 2D captions, as they provide critical, fine-grained object-level attributes that refine the referential grounding.
+We see that SORT3D is able to explainably resolve complex view-dependent relations with multiple anchors and complex semantic descriptions (Figure 4-a), while also providing explainable model failure points by analyzing its chain of thought reasoning (Figure 4-b).
+On IRef-VLA, our method surpasses other methods on Hard statements by a large margin.
+IRef-VLA contains a large set of statements using size and color descriptions when referring to objects, which our method effectively grounds by utilizing object filtering, semantic attributes captured by captions, and the spatial reasoning toolbox to identify target objects with multiple distractors in the scene.
+Transcrib3D, when used in its zero-shot formulation with principles and not fine-tuned on the dataset’s hard statements [6], fails to generalize well despite using the same LLM backbone.
+The results on IRef-VLA strongly support the benefits of incorporating 2D captions, as they provide critical, fine-grained object-level attributes that refine the referential grounding.
 Figure 4: Correct (a) and incorrect (b) grounding examples.
@@
 In the bottom right, the model fails at pragmatics, picking out the rightmost pillow, instead of recognizing that the sentence implies choosing a pillow on the bed.
-TABLE I: Performance on the Nr3D dataset from the ReferIt3D benchmark.
+TABLE II: Performance on the Nr3D dataset from the ReferIt3D benchmark.
 Asterisks (*) indicate results reported from the paper directly.
-“+skip” indicates exclusion of statements answered incorrectly by human annotators.
 “View Dep.” and “View Ind.” stand for view-dependent and view-independent respectively.
@@
 SceneVerse-GPS* [19] 64.9 56.9 67.9
-Fine-Tuned Methods
-Transcrib3D* [6] (GPT-4+skip) 70.2 60.1 75.4
-Transcrib3D (GPT-4o) 58.5 54.0 60.6
-Transcrib3D (Mistral) 58.5 52.4 61.3
-Transcrib3D (GPT-4o+skip) 65.6 63.3 66.7
-Transcrib3D (Mistral+skip) 63.8 57.1 66.7
 Zero-Shot Methods
-ZSVG3D [20] 39.0 36.8 40.0
-VLM-Grounder [11] 48.0 45.8 49.4
-CSVG [9] 59.2 53.0 62.5
-Ours (Mistral) 56.5$\pm$0.5 57.6$\pm$3.6 56.2$\pm$2.3
-Ours (GPT-4o) 60.5$\pm$1.0 56.6$\pm$3.3 62.3$\pm$2.4
-Ours (GPT-4o+skip) 62.0$\pm$1.2 56.6$\pm$0.0 64.3$\pm$1.7
-Ours (Mistral+skip) 61.6$\pm$0.3 59.4$\pm$0.9 62.6$\pm$0.9
-TABLE II: Performance on the Sr3D dataset from the ReferIt3D benchmark. Asterisks (*) indicate results reported from the paper directly.
+ZSVG3D* [21] 39.0 36.8 40.0
+VLM-Grounder* [11] 48.0 45.8 49.4
+CSVG* [9] 59.2 53.0 62.5
+Transcrib3D* [6] (GPT-4) 70.2 60.1 75.4
+Transcrib3D (GPT-4o) 65.6 63.3 66.7
+Transcrib3D (Mistral) 63.8 57.1 66.7
+[…]
+[…]
+TABLE III: Performance on the Sr3D dataset from the ReferIt3D benchmark. Asterisks (*) indicate results reported from the paper directly.
 Sr3D
@@
 SceneVerse-GPS* [19] 77.5 62.8 78.2
-Fine-Tuned Methods
-Transcrib3D* [6] (GPT-4+skip) 98.4 98.2 98.4
+Zero-Shot Methods
+Transcrib3D* [6] (GPT-4) 98.4 98.2 98.4
 Transcrib3D (GPT-4o) 96.5 88.9 96.9
 Transcrib3D (Mistral) 96.0 77.8 96.9
-Zero-Shot Methods
-Ours (Mistral) 92.0$\pm$0.7 90.9$\pm$0.0 92.2$\pm$0.8
-Ours (GPT-4o) 92.0$\pm$0.0 95.5$\pm$0.0 91.6$\pm$0.0
-TABLE III: Grounding accuracy on VLA-3D test subset
-VLA3D
+[…]
+[…]
+TABLE IV: Grounding accuracy on IRef-VLA test subset
+IRef-VLA
 Method Overall Easy Hard
@@
 Transcrib3D (GPT-4o) 77.5 82.5 57.5
-Ours (Mistral) 69.0$\pm$0.7 70.0$\pm$0.8 65.0$\pm$0.0
-Ours (GPT-4o) 71.8$\pm$1.8 71.0$\pm$1.3 75.0$\pm$3.5
+[…]
+[…]
 V-B Ablation of Captioning Module
@@
 We hypothesize that these finer descriptions of object attributes will help the model in disambiguating objects when given free-form referential statements.
-We test this hypothesis through an ablation study using GPT-4o with both our method and Transcrib3D on Nr3D, shown in Table IV.
+We test this hypothesis through an ablation study using GPT-4o with both our method and Transcrib3D on Nr3D, shown in Table V.
 In both Transcrib3D and our method, we observe consistent and significant improvements across all statements types.
@@
 Many referential statements rely on subtle distinctions—such as color, shape, texture, or affordance—that traditional 3D models often miss.
-V-C Real-World Validation
-To validate SORT3D’s pipeline in the real-world, we implemented the full system on a differential-drive ground vehicle equipped with a 360-degree camera, a Velodyne LiDAR sensor, an onboard Intel NUC for low level autonomy, and a desktop with an RTX 4090.
-The system starts with an empty map and continuously integrates real-time RGB-D input to perform panoptic segmentation.
-This enables the robot to simultaneously build a map, detect objects, and assign open-vocabulary labels, allowing it to navigate and understand its environment dynamically.
-As the robot explores, it captures images of objects and our pipeline selects image crops and generates open-vocabulary captions, then responds to grounding language queries whenever they are typed in real-time, triggering the rest of the SORT3D pipeline.
-We validate our system in a previously unseen environment, demonstrating its ability to handle complex statements including compound statements and open-vocabulary object-centric navigation commands, as shown in Fig 6.
-Figure 5: An overhead view of the real world kitchen environment and the differential-drive robot used in the experiment.
-Figure 6: Real-world examples of SORT3D referential grounding and navigation on compound relational statements and multi-object navigation 3D.
-The top row shows the egocentric camera view of the robot and the bottom row shows the path taken in a map of the scene.
-The target objects are highlighted in green and the anchor objects blue.
-VI LIMITATIONS AND FUTURE WORK
-We acknowledge that limitations exist within in our approach.
-First, our system relies on internet access to call online APIs, which is a reasonable assumption for indoor environments like homes but may not hold for generalizable robot navigation systems.
-To address this limitation, the modular design of our pipeline allows for the LLMs to be easily replaced by smaller local models.
-We also recognize that our evaluation set is limited.
-There are few existing benchmarks designed to rigorously test online 3D referential grounding with diverse, attribute-rich natural language.
-This constraint makes it challenging to fully assess the generalizability of our approach across different environments or while deployed on a system.
-Evaluating our method in a simulated environment with real-time interactions would provide deeper insights into its effectiveness and adaptability as a system.
-TABLE IV: Grounding accuracy with and without captions on Nr3D
+TABLE V: Grounding accuracy with and without captions on Nr3D
 Nr3D
@@
 Ours (GPT-4o + captions) 60.5 ($\uparrow$ 11.0%) 64.2 ($\uparrow$ 8.4%) 55.0 ($\uparrow$ 5.3%) 56.6 ($\uparrow$ 11.6%) 62.3 ($\uparrow$ 10.9%)
+V-C Real-World Validation
+To validate SORT3D’s pipeline in the real-world, we implement the full system on two autonomous mobile robots: a mecanum-wheeled robot and a differential drive robot with a wheelchair base, each equipped with 360-degree cameras, respective Livox and Velodyne LiDARS, onboard Intel NUCs for low level autonomy, and RTX 4090s with 16GB and 24GB of VRAM respectively.
+We validate our system in two previously unseen indoor environments: a student lounge (Figure 6, and a university corridor (Figure 5).
+We attempt two different types of navigational queries, which target the system’s ability to ground statements that employ both spatial references and object semantic attributes.
+Before issuing a query, we navigate each robot around the scene to build a semantic map after prompting the open vocabulary detector with the names and synonyms of objects in the scene (shown in RViz in the bottom of Figures 6 and 5), and collect image crops for each detected object.
+As an implementation detail, captions are batch-generated only after a user query is typed into the system to speed up semantic mapping and decrease overall power consumption.
+RViz visualizations of the instance level semantic maps, the objects and corresponding waypoints chosen by the grounding model for each statement, and pictures of each platform navigating through its environment are shown in Figures 6 and 5.
+In each statement we test, SORT3D successfully grounds one or more referenced objects, demonstrating the versatility of our approach for grounding complex expressions involving spatial references and semantic attributes in previously unseen scenes.
+Further experiments on more environments and types of statements are found in our accompanying repository and videos.
+Figure 5: Navigation on the mecanum robot in the university corridor given the statement “Go to the table next to the bookshelf, then to the chair next to the plant.” The system successfully navigates to both spatially referenced objects.
+Figure 6: Navigation on the mecanum robot in the student lounge given the statement “I want to play a board game, fetch me one from the shelf.” The system successfully navigates to the shelf with the board game on it by referring to semantic caption descriptions.
+VI Limitations and Future Work
+We acknowledge that limitations exist within our approach.
+First, our system relies on internet access to call online APIs, which is a reasonable assumption for indoor environments like homes and offices, but may not hold for other environments where generalizable robot systems may need to be deployed.
+To address this limitation, the modular design of our pipeline allows for the LLMs to be easily replaced by smaller local models.
+We also recognize that our evaluation set is limited.
+There are few existing benchmarks designed to rigorously test online 3D referential grounding with diverse, attribute-rich natural language.
+This constraint makes it challenging to fully assess the generalizability of our approach across different environments or while deployed on a system.
+Evaluating our method in a simulated environment with real-time interactions would provide deeper insights into its effectiveness and adaptability as a system.
+It would also allow us to investigate more practical settings such as multi-turn grounding scenarios and failure correction.
 VII Conclusion